OpenStudy (anonymous):

Find the equation of a line parallel to the given line: y = 2x + 4. y = -2x + 4 y = 2x y = 1/2x + 4 y = -1/2x -4

OpenStudy (anonymous):

parallel means same slope, therefore y = 2x is the correct answer.

OpenStudy (theviper):

\[\frac{a_{1}}{a_{2}}=\frac{b_{1}}{b_{2}} \neq \frac{c_{1}}{c_2}\]

OpenStudy (anonymous):

Sorry? I'm not sure what you meant by that ^